OpenStudy (anonymous):

Now let's look at some real numbers. For a big transmitting amplifier the output resistance may be R1=1000.0Ω. A typical antenna has a radiation resistance of R2=50.0Ω. Consider an AM broadcast transmitter at f=840.0kHz. In the spaces provided below, write the numerical values of the capacitance (in picoFarads) and inductance (in microHenrys) for match.

OpenStudy (anonymous):

w = 2*pi*f Cmatch = 1/(w*sqrt(R1R2 – R2^2)) Cmatch 869.34 Lmatch = sqrt(R1R2 – R2^2)/w Lmatch 41.29

OpenStudy (rva.raghav):

abhay give some books to prefer for multivibrators